LaurynCarroll:

Shelby bought a painting for dollars. She was charged a 7% sales tax for the painting. Which expression correctly shows the total price that Shelby paid for the painting?

Nightmoon:

Shelby bought a painting for x amount of dollars Sales Tax = 7% Sales Tax= 7% of x Sales Tax = 0.07 * x Sales Tax = 0.07x Total price = Selling Price + Sales Tax Total Price = x + 0.07x Ending result: She Paid 1.07x for the painting -